A football is thrown with an initial velocity of 43 mph at an angle of 35 degrees. How many yards down the field does the ball land?
1
Expert's answer
2018-09-19T15:39:08-0400
1 mph=0.44704 m/s
R=v^2/g sin⁡〖2(θ)〗
R=〖(0.44704∙43)〗^2/9.8 sin⁡70=35.43 m=35.43(1.09361)yards=39 yards.